### v2

# L-FUSION: Laplacian Fetal Ultrasound Segmentation & Uncertainty Estimation

Abstract
Accurate analysis of prenatal ultrasound (US) is essential for early detection of developmental anomalies. However, operator dependency and technical limitations (e.g. intrinsic artefacts and effects, setting errors) can complicate image interpretation and the assessment of diagnostic uncertainty. We present L-FUSION (Laplacian Fetal US Segmentation with Integrated FoundatiON models), a framework that integrates uncertainty quantification through unsupervised, normative learning and large-scale foundation models for robust segmentation of fetal structures in normal and pathological scans. We propose to utilise the aleatoric logit distributions of Stochastic Segmentation Networks and Laplace approximations with fast Hessian estimations to estimate epistemic uncertainty only from the segmentation head. This enables us to achieve reliable abnormality quantification for instant diagnostic feedback. Combined with an integrated Dropout component, L-FUSION enables reliable differentiation of lesions from normal fetal anatomy with enhanced uncertainty maps and segmentation counterfactuals in US imaging. It improves epistemic and aleatoric uncertainty interpretation and removes the need for manual disease-labelling. Evaluations across multiple datasets show that L-FUSION achieves superior segmentation accuracy and consistent uncertainty quantification, supporting on-site decision-making and offering a scalable solution for advancing fetal ultrasound analysis in clinical settings.
Keywords:
Fetal Ultrasound Foundation Models Uncertainty Quantification Segmentation
1 Introduction
Fetal ultrasound imaging is a crucial tool for monitoring fetal development and facilitating the early detection of potential abnormalities, enabling timely interventions that support optimal care. Detection rates for abnormalities have significantly improved, rising from $15-59~{}\%$ in the 2000s to $30-88~{}\%$ in recent years, depending on the type of lesion [6]. However, ultrasound images are noisy and contain shadows, variability and ambiguities in anatomy, creating challenges for clinicians to precisely identify subtle, early-stage, or rare abnormalities. Therefore, computerised support using machine learning models is expected to improve outcomes. Such models can mitigate operator dependence by providing consistent, objective assessments, and for selected developmental problems, their detection rates are often on par with human observations [5, 21]. However, these models, when designed to provide predictions without any confidence bounds, may lack the ability to convey their certainty to the operator. In a safety-critical domain such as healthcare, modelling uncertainty quantification can improve diagnostic accuracy and also assist clinicians in focusing their attention on critical regions by effectively flagging areas of high uncertainty. Two main constraints limit the application of those models: Firstly, methods for uncertainty estimation suffer from high computational costs [26], especially, with an increasing number of imaging dimensions. Laplacian uncertainty approximation emerges as a particularly powerful framework, enabling fine-grained distinctions between anatomical variations seen in healthy and pathological cases. Furthermore, by incorporating faster Hessian approximations the framework keeps computational requirements tractable and speeds up compute time. Secondly, the main challenge for effective probabilistic approaches and efficient uncertainty estimation has been their requirement for large datasets. Epistemic uncertainty captures gaps in model knowledge, while aleatoric uncertainty accounts for noise and inherent variability in the data [10]. Epistemic uncertainty can be decreased by increasing the amount of data but is required to capture situations not encountered in the training set. Especially, in safety-critical systems, epistemic uncertainty is used to detect situations which have never been seen by the model before [13]. However, data scarcity, along with challenges in robustness and generalization, continues to be critical issues in ultrasound imaging. Integrating pretrained large-scale foundation models [11] in our Laplacian framework can address these challenges by leveraging their ability to generalise across diverse datasets while minimising the need for extensive retraining. These models are believed to transfer knowledge from large-scale training data to specialised domains, reducing training costs while significantly improving diagnostic performance. The benefits are twofold: first, they improve detection rates for subtle and complex conditions, and second, they ensure adaptability to variable imaging environments and limited data scenarios. When combined with advanced uncertainty frameworks like Laplacian approximations, these systems are uniquely equipped to tackle the statistical and computational requirements of uncertainty estimation in challenging datasets.
Contributions: We propose L-FUSION, a novel approach that integrates Laplacian uncertainty approximation with embeddings from a foundation model to deliver accurate segmentations and precise uncertainty quantification for fetal ultrasound. The key contributions of this work are:
(1) We utilise a deterministic ultrasound foundation model and facilitate aleatoric and epistemic uncertainty quantification through Laplacian approximation of weight posteriors of a segmentation head with Gaussian distributions. Our method achieves superior segmentation performance in both binary and multi-class applications, while enhancing unsupervised out-of-distribution detection by incorporating additional uncertainty quantification and segmentation counterfactual through a dropout component.
(2) By restricting the computationally intensive Hessian calculation for the Laplace approximation to the segmentation head, we enhance our method by integrating a fast Hessian approximation with an efficient, versatile encoder. This results in a highly efficient and generalisable solution for universal ultrasound imaging tasks in clinical applications.
(3) We introduce a 3D extension of our method that combines a deterministic encoder with a Laplacian segmentation head for ultrasound clips. For comparison, we developed a 3D Laplacian segmentation model, as no universal ultrasound video foundation model currently exists for optimal benchmarking.
Related Work: Despite its widespread clinical use, automated segmentation of ultrasound imaging lags behind other imaging modalities (e.g. lack of standardised performance measures tailored for US, smaller datasets, high inter- and intra-expert variability) [19]. However, there is a growing emphasis on better validation, including comparisons across methods and evaluation on larger, standardised datasets, to support broader clinical adoption [19]. Recent research in ultrasound segmentation has shifted towards more systematic use of priors, such as geometric, temporal, and imaging physics constraints, to improve segmentation accuracy [18, 23, 14, 8]. Foundation models, trained on large-scale natural image datasets, have been adapted for various medical imaging modalities, such as X-ray, computer tomography, magnetic resonance imaging, and ultrasound. These models learn modality-specific features, but their ability to generalise across modalities remains limited [25]. In ultrasound imaging, two notable foundation models have recently emerged to address challenges specific to this modality. The Multi-Organ FOundation (MOFO) [2] model seeks to overcome data scarcity and the limitations of single-organ models by segmenting images across multiple organs. By extracting organ-invariant representations and incorporating anatomical priors, MOFO enhances segmentation accuracy by exploiting relationships between anatomical structures from different regions. Evaluated on a multi-organ database, MOFO outperforms single-organ segmentation methods in various metrics. With a similar aim, the Universal ultrasound Foundation Model (USFM) [11] tackles generalisation issues across tasks and organs by training on a large-scale multi-organ, multi-centre, and multi-device database. USFM uses self-supervised pre-training with a spatial-frequency dual-masked image modelling method to extract robust features from low-quality images. Foundation models extract rich representations, but they remain indifferent to uncertainty. Bayesian modelling provides a probabilistic approach to segmentation, quantifying both epistemic and aleatoric uncertainty [15, 1, 3, 7]. Laplacian approximations can be employed within a Bayesian framework to estimate epistemic uncertainty by approximating the posterior distribution. However, the direct computation of the Hessian is infeasible, as it scales quadratically with the number of model parameters [4]. Recent work [24] combines the aleatoric logit distribution from Stochastic Segmentation Networks [17] with Laplace approximations for epistemic uncertainty quantification.
Figure 1: Foundation model (Encoder) with Laplacian segmentation head $L(f)$ and a dropout unit. (a) $L(f)$ decoding the embeddings $f$ from the Foundation model $\bar{F}(x)$ and a Dropout U-Net $G(x)$. (b) $L(f)$ and Dropout Segmentation Head $G(f)$ decoding the embeddings $f=\bar{F}(x)$. For both approaches, the segmentations $S_{i}(f)$ are predicted by $L(f)$. We enhance the uncertainty maps $U$ with OOD Calibration by including the variance and total entropy of logit differences.
2 Method
As illustrated in Fig. 1 the L-FUSION framework integrates three key components: (1) USFM [11] as the encoder, (2) a Laplacian segmentation head and (3) a Dropout unit for combined segmentation and uncertainty quantification with enhanced uncertainty maps.
Laplacian Component Given an input fetal ultrasound image $x$, the pretrained foundation model processes $x$ to generate embeddings $f$. All weights remain deterministic and frozen (i.e., no further fine-tuning or randomisation), which significantly reduces the complexity and computational cost typically involved in training deep neural networks. The embeddings $f$ are then passed to the Laplacian segmentation head $L$, which produces a normal distribution of logits $\eta_{L}|f\sim\mathcal{N}(\mu(f,\theta_{1}),\varSigma(f,\theta_{2}))$. $\theta_{1}$ and $\theta_{1}$ parametrise the distribution of segmentation head networks. Next, we sample the segmentation mask $S_{i}$ from the Bernoulli distribution parametrised by the sigmoid of the respective logit $\eta_{L,i}$, for $i=1,...,n$ with $n$ samples per input. For aleatoric uncertainty, we Monte Carlo-sample $m$ segmentation networks $L_{j}(f)$ for $j=1,...,m$ from the Laplace approximation $q(\theta^{*})$ for an embedding $f=\bar{F}(x)$. Each sampled segmentation network predicts one logit distribution $p(\eta|f,\theta)$ for each embedding $f$. Finally, we get the mean prediction map $\hat{S}_{L}$ and the uncertainty maps $U_{L,k}$, for $k=1,...,K$ with K uncertainty measures, expressed as $\hat{S}_{L},U_{L,k}=L(\bar{F}(x))$, with the frozen encoder $\bar{F}$. $U_{L,k}^{E}$ reflects epistemic uncertainty by measuring the mutual information $\mathcal{I}(p(S,\theta^{*}|f,\mathcal{F}))$, with the entirety of embeddings $\mathcal{F}$, indicating the model’s confidence and highlighting areas with less reliable predictions. $U_{L,k}^{A}$ captures aleatoric uncertainty by calculating the expected entropy $\mathcal{H}(p(S|f,\theta^{*})$, which results from data noise such as poor imaging quality or natural variability in fetal anatomy. Given the low-rank matrix of the Hessian $P$, $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D)$ defines the low-rank parametrisation of the covariance matrix [17] and $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D)$ the fast Hessian approximations [24]. Both allow faster computation of the Laplace approximation.
Theorem 2.1 (Sufficiency of Pretrained Embeddings for Segmentation and Uncertainty Quantification)
Let $x\in\mathcal{X}$ be an input image, and let $\bar{F}:\mathcal{X}\to\mathcal{F}$ be a deterministic, pretrained foundation model mapping $x$ to a feature embedding $f=\bar{F}(x)$. Consider a Laplacian segmentation head $L$ that models the posterior segmentation probability as $p(S|f,\theta)=\prod_{i}\text{Bernoulli}(s_{i}|\sigma(\eta_{L,i})),$ where the logit distribution follows the Laplacian approximation, $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D).$ If $\bar{F}$ preserves segmentation-relevant information, i.e., $I(S;x)\newline =I(S;f),$ then $f$ is a sufficient statistic for segmentation uncertainty estimation.
Proof
Since $\bar{F}$ is deterministic, the segmentation posterior simplifies to $p(S|x)=p(S|f,\theta),$ which holds if $p(f|x)$ is a delta distribution centered at $\bar{F}(x)$, meaning $p(f|x)=\delta(f-\bar{F}(x)).$ By the Data Processing Inequality [20], $I(S;x)\geq I(S;f),$ and since $f$ is a deterministic function of $x$, we get $I(S;x)=I(S;f),$ ensuring that $f$ is a sufficient statistic for segmentation. For uncertainty estimation, the Laplacian approximation assumes a Gaussian weight posterior, $q(\theta^{*})=\mathcal{N}(\theta^{*}|\theta_{\text{MAP}},\mathbf{H}^{*-1}),$ which propagates uncertainty into the predictive variance: $\operatorname{Var}(S|f)=J\mathbf{H}^{*-1}J^{T}+\sigma^{2}(f)I.$ Here, $J\mathbf{H}^{*-1}J^{T}$ captures epistemic uncertainty, and $\sigma^{2}(f)$ captures aleatoric uncertainty. Since both uncertainties depend only on $f$, it follows that $f$ is sufficient for segmentation and uncertainty quantification.
Dropout Component Additionally, we use a Dropout unit, a Dropout U-Net $G(x)$ (Fig. 1 a) or a Dropout Segmentation Head $G(f)$ (Fig. 1 b) taking the embeddings from the Foundation Model encoder as inputs. From the Dropout unit, we take the logit distribution $\eta_{MC}$ for sampling logits and epistemic uncertainty measures. For uncertainty quantification, we propose an OOD-Calibration mechanism based on logit variance $U_{V}=Var(\eta_{q})$ (Intra-Model) or logit difference variance between $U_{V}=Var(\left(\frac{1}{q}\sum_{q=1}^{Q}\eta_{q}\right)$ for $Q$ models (Inter-Model), and scaled by OOD measure from the same model (Intra-Model) or other model(s) (Inter-Model) for improving OOD detection AUC. For OOD calibration, we generate the uncertainty maps by including the total entropy $\mathcal{H}(U_{V})$ through its Hadamard product with $U_{V}$ which scales and limits the uncertainty maps to only areas of total entropy larger than zero in $U_{V}$.
3 Evaluation
Our fetal heart dataset consists of US scans in four-chamber view (4CH), with training data ($n=428$) comprising only normal fetal heart samples, we include $6$ label classes (all chambers, heart and thorax). As OOD samples we test on an example of fetal Congenital Heart Disease (CHD), AtrioVentricular Septal Defect (AVSD) in 4CH. The normal test set includes $92$ normal and $193$ AVSD samples. The public HC18 dataset [9] consists of 1,334 2D US images of fetal standard planes used to measure head circumference (HC), a key metric for assessing gestational age and fetal growth. It includes 999 training images with expert annotations and measurements, and 335 test images without annotations. Images are $800\times 540$ pixels with pixel sizes ranging from 0.052 to 0.326 mm, provided in accompanying metadata files, accessible via Zenodo. As OOD samples we use synthetically generated local deformations on the skull resulting in abnormal skull shape.
Implementation and Training: We extended the $nnj$ library [16], which is used for implementing layers for diagonal backpropagation and fast hessian approximations, by implementing a BatchNorm to reduce exploding gradients and add support for higher-dimensional data (Conv3d, Upsample3d, Batchnorm3d). We resize images to $256\times 256$ for U-Net-based architectures and $224\times 224$ for the UMSF encoder[11]. All evaluated U-Nets have the same architecture besides the Dropout U-Net [12] which has Dropout-layers after each convolutional block. We augment training images with linear transformations and random crop up to $80\%$ of the original size. We align fetal heart scans with an iterative approach, a 2D variant of [22]. All models were trained on an NVIDIA A100 $80$ GB with Early stopping, the number of epochs limited to $250$ for the HC18 dataset and to $400$ (2D) and $600$ (Clip) for the fetal datasets at maximum.
Metrics: Segmentation performance is evaluated using Dice Similarity Coefficient (DSC), Hausdorff Distance (HD), and, Absolute Difference (AD) for HC18 according to the related challenge or Intersection over Union (IoU) for the fetal heart datasets, to quantify overlap and boundary alignment. For out-of-distribution (OOD) performance, we follow [24] and report the Area Under the Receiver Operating Characteristic Curve (AUROC/AUC) across epistemic components, specifically Mutual Information (MI), Expected Pairwise KL (EP), Pixel Variance (PV), and aleatoric uncertainty with Expected Entropy (EE).
Table 1: Results on Head Circumference Dataset HC18; DSC [%] - Dice Similarity Coefficient, HD [mm] - Hausdorff Distance, AD [mm] - Absolute Difference; 1st-ranked, 2nd-ranked.
Mean 25-Percentile 75-Percentile 1. Trimester
Model DSC $\uparrow$ HD $\downarrow$ AD $\downarrow$ DSC $\uparrow$ HD $\downarrow$ AD $\downarrow$ DSC $\uparrow$ HD $\downarrow$ AD $\downarrow$ DSC $\uparrow$ HD $\downarrow$ AD $\downarrow$
SSN [17] $87.6_{\pm 15.0}$ $8.4_{\pm 9.1}$ $16.7_{\pm 22.5}$ 85.7 1.9 2.9 96.6 10.9 19.9 $73.2_{\pm 23.5}$ $13.7_{\pm 9.6}$ $23.3_{\pm 29.6}$
Drop. U-Net [12] $\underline{94.3_{\pm 8.0}}$ $\underline{4.1_{\pm 4.3}}$ $\underline{9.2_{\pm 11.3}}$ 93.9 1.8 3.2 97.2 4.9 10.5 $\underline{87.7_{\pm 17.6}}$ $\underline{4.2_{\pm 7.3}}$ $\underline{10.1_{\pm 21.0}}$
Lap. U-Net [24] $85.4_{\pm 16.9}$ $9.9_{\pm 11.7}$ $21.9_{\pm 31.6}$ 82.7 2.7 4.3 96.3 11.6 25.2 $70.3_{\pm 22.2}$ $11.3_{\pm 10.6}$ $29.7_{\pm 30.1}$
L-FUSION (ours) $\mathbf{95.5_{\pm 8.0}}$ $\mathbf{2.7_{\pm 3.0}}$ $\mathbf{4.7_{\pm 6.2}}$ $\mathbf{95.9}$ $\mathbf{1.3}$ $\mathbf{1.4}$ $\mathbf{97.7}$ $\mathbf{2.9}$ $\mathbf{5.6}$ $\mathbf{89.0_{\pm 17.7}}$ $\mathbf{3.4_{\pm 5.6}}$ $\mathbf{6.4_{\pm 10.8}}$
Quantitive Results: L-FUSION outperforms all baselines on HC18, Tab. 1. It achieves the highest DSC ($95.5~{}\%$) and the lowest HD ($2.7~{}mm$) and AD ($4.7~{}mm$), surpassing Dropout U-Net by $1.3~{}\%$ in DSC and reducing HD by $34.1~{}\%$. Performance gains are consistent across percentiles and the first trimester, where L-FUSION improves DSC by $1.5~{}\%$ over Dropout U-Net while reducing AD by $36.6~{}\%$. Results demonstrate the superiority of L-FUSION in head circumference segmentation. For fetal heart ultrasound (Tab. 2), L-FUSION sets new benchmarks. In the 4CH plane without alignment, it outperforms the best baseline by $12.8\%$ in DSC and reduces HD by $41.6\%$. Under thorax- and heart-aligned conditions, it leads by $9.5\%$ and $0.9\%$ in DSC, respectively, while consistently achieving the lowest HD. L-FUSION also cuts training time by up to $40\%$, with greater savings under stronger alignment, affirming its state-of-the-art status in fetal heart and head circumference segmentation.
Table 2: Segmentation Performance on Fetal Hearts over all class labels; Plane - Fetal heart US standard plane; A. - Aligned to Thorax (T) or Heart (H); DSC [%] - Dice Similarity Coefficient, IoU [%] - Intersection over Union, HD [pixel] - Hausdorff Distance; 1st-ranked, 2nd-ranked.
SSN [17] Dropout U-Net [12] Laplace U-Net    [24] L-FUSION (ours)
Plane A. DSC $\uparrow$ IoU $\uparrow$ HD $\downarrow$ DSC $\uparrow$ IoU $\uparrow$ HD $\downarrow$ DSC $\uparrow$ IoU $\uparrow$ HD $\downarrow$ DSC $\uparrow$ IoU $\uparrow$ HD $\downarrow$
2D 4CH - $40.6_{\pm 10.4}$ $26.1_{\pm 8.8}$ $\underline{44.0_{\pm 7.1}}$ $43.0_{\pm 9.0}$ $27.9_{\pm 7.7}$ $62.3_{\pm 16.9}$ $\underline{43.1_{\pm 10.1}}$ $\underline{28.0_{\pm 8.9}}$ $49.8_{\pm 9.0}$ $\mathbf{55.9_{\pm 8.6}}$ $\mathbf{39.3_{\pm 9.1}}$ $\mathbf{29.1_{\pm 4.7}}$
4CH T $47.7_{\pm 11.7}$ $32.1_{\pm 10.4}$ $63.4_{\pm 12.7}$ $60.2_{\pm 5.9}$ $43.3_{\pm 6.3}$ $89.5_{\pm 42.3}$ $\underline{61.2_{\pm 7.6}}$ $\underline{44.5_{\pm 7.8}}$ $\underline{37.0_{\pm 10.7}}$ $\mathbf{67.0_{\pm 7.9}}$ $\mathbf{50.9_{\pm 8.5}}$ $\mathbf{36.9_{\pm 15.1}}$
4CH H $52.4_{\pm 12.2}$ $36.5_{\pm 12.2}$ $\underline{56.1_{\pm 16.7}}$ $\underline{65.0_{\pm 6.7}}$ $\underline{48.5_{\pm 7.8}}$ $68.1_{\pm 26.0}$ $55.9_{\pm 9.3}$ $39.5_{\pm 9.9}$ $56.8_{\pm 14.3}$ $\mathbf{65.6_{\pm 6.8}}$ $\mathbf{49.2_{\pm 7.9}}$ $\mathbf{35.0_{\pm 10.7}}$
Clip 4CH $46.1_{\pm 12.4}$ $30.9_{\pm 11.9}$ $\underline{23.7_{\pm 4.9}}$ $33.7_{\pm 12.9}$ $21.1_{\pm 10.3}$ $83.8_{\pm 32.4}$ $\underline{53.7_{\pm 7.8}}$ $\underline{37.1_{\pm 7.2}}$ $23.8_{\pm 5.0}$ $\mathbf{66.8_{\pm 8.1}}$ $\mathbf{51.3_{\pm 7.5}}$ $\mathbf{23.4_{\pm 5.2}}$
4CH H $\underline{68.1_{\pm 5.9}}$ $\underline{51.9_{\pm 6.6}}$ $\mathbf{20.8_{\pm 3.7}}$ $67.4_{\pm 6.0}$ $51.1_{\pm 6.5}$ $32.2_{\pm 15.6}$ $\mathbf{71.2_{\pm 5.3}}$ $\mathbf{55.5_{\pm 6.3}}$ $\underline{21.3_{\pm 3.9}}$ $65.6_{\pm 6.4}$ $49.1_{\pm 6.9}$ $22.3_{\pm 3.9}$
Tab. 3 compares the best-performing individual OOD measures for epistemic and aleatoric uncertainty against OOD-Calibration, which scales predicted logit variance ($n=50$) to enhance OOD detection AUC. Applied within the same model (Intra-Model) or across collaborating models (Inter-Model), OOD-Calibration significantly improves AUC. On unaligned 4CH and HC18 datasets, Inter-Model calibration increased AUC by $19\%$ (HC18) and $20\%$ (4CH). For the aligned 4CH dataset, Intra- and Inter-Model calibration performed equally, yielding AUC gains of $3-9\%$.
Table 3: AUC for OOD Detection with OOD-Calibration (using the best-scoring OOD measure for AUC: Expected Entropy, Expected Pairwise Kullback-Leibler, Pixel Variance, Mutal Information); A. - Aligned to Thorax (T) or Heart (H); D. - Dropout, U. - U-Net, L. - Laplace, B - UMSF $+$ D. Segmentation Head, A - UMSF $+$ D.U.; HC18 - Synthetic HC18 datset with Clips;1st-ranked, 2nd-ranked.
OOD - Calibration
Indiv. OOD Measures Intra-Model Inter-Model
Data A. D.U.[12] L.U.[24] L.[11] D.[11] D.U.[12] L.U.[24] L.[11] D.[11] B A
2D HC18 $0.74^{\dagger}$ $0.57^{\dagger}$ $0.70^{\dagger}$ $0.72^{\circ}$ $0.84^{\ast}$ $0.52^{\dagger}$ $0.71^{\dagger}$ $0.74^{\dagger}$ $\mathbf{1.00}^{\ast}$ $\mathbf{1.00}^{\ast}$
4CH $0.61^{\circ}$ $0.56^{\circ}$ $0.57^{\dagger}$ $0.61^{\circ}$ $0.61^{\circ}$ $0.51^{\circ}$ $0.54^{\ast}$ $0.67^{\ast}$ $\mathbf{0.84}^{\ast}$ $\underline{0.73}^{\ast}$
4CH T $0.61^{\circ}$ $0.54^{\circ}$ $0.56^{\circ}$ $0.56^{\circ}$ $\mathbf{0.89}^{\ast}$ $0.53^{\ast}$ $0.54^{\ast}$ $0.72^{\ast}$ $0.64^{\ast}$ $\underline{0.82}^{\dagger}$
4CH H $0.64^{\diamond}$ $0.59^{\dagger}$ $0.55^{\ast}$ $0.56^{\ast}$ $\mathbf{0.85}^{\circ}$ $0.54^{\ast}$ $0.51^{\circ}$ $0.64^{\circ}$ $0.71^{\ast}$ $\underline{0.82}^{\dagger}$
Clip HC18 $0.73^{\circ}$ $0.57^{\dagger}$ $0.69^{\dagger}$ $0.70^{\circ}$ $0.82^{\ast}$ $0.53^{\dagger}$ $0.73^{\circ}$ $0.75^{\dagger}$ $\underline{0.98^{\ast}}$ $\mathbf{0.99^{\ast}}$
4CH $\mathbf{0.88^{\dagger}}$ $0.77^{\dagger}$ $0.71^{\ast}$ $\underline{0.81^{\diamond}}$ $\underline{0.81^{\circ}}$ $0.58^{\dagger}$ $0.55^{\circ}$ $\mathbf{0.88^{\circ}}$ $\mathbf{0.88^{\circ}}$ $\underline{0.81^{\circ}}$
4CH H $\underline{0.67^{\dagger}}$ $\underline{0.67^{\circ}}$ $0.66^{\circ}$ $0.62^{\ast}$ $\underline{0.67^{\circ}}$ $0.59^{\circ}$ $\underline{0.67^{\circ}}$ $\mathbf{0.73^{\ast}}$ $\mathbf{0.73^{\ast}}$ $\underline{0.67^{\dagger}}$
Figure 2: Segmentation and uncertainty maps predicted by L-FUSION for Normal and OOD (AVSD) cases in four-chamber view (4-CH) Heart-aligned.
Qualitative Results: Dropout-based uncertainty maps often produce diffuse estimates, see Fig. 1, tending to over-segmentation and missing small details, hence, failing to model fine-grained epistemic uncertainty. The foundation model with a Laplacian segmentation head improves segmentation accuracy and benefits from fine-grained uncertainty estimates but shows high global uncertainty. Furthermore, L-FUSION’s variance maps of logit differences enable counterfactual analysis and improve visual uncertainty maps from both intra- and inter-model variations, see Fig. 2. This provides a more reliable uncertainty quantification, crucial for applications requiring transparent model confidence.
Discussion: Our approach enables efficient segmentation with fine-grained uncertainty quantification benefiting from two uncertainty-modelling approaches, Laplacian approximation and Monte-Carlo Dropout. With OOD-Calibration, we pushed OOD detection AUC up to $+20\%$. However, performance may decline when domain-specific features are essential or data lies on complex, nonlinear manifolds. Uncertainty quantification also depends on embedding quality—loss of local details or high noise can impair segmentation.
4 Conclusion
Ultrasound segmentation has advanced but still lags behind CT and MRI in uncertainty quantification and consistency[19]. Foundation models [2, 11] improve accuracy but often neglect uncertainty quantification, crucial for clinical use. By combining a foundation model encoder with a Laplacian segmentation head with fast Hessian approximation, a Dropout unit and OOD-Calibration, we reduce complexity while achieving superior performance. Our approach enhances accuracy and generates localised uncertainty maps, improving reliability and clinical applicability over deterministic and Dropout U-Nets.
Acknowledgements: We thank the volunteers and sonographers at St. Thomas’ Hospital London. We gratefully acknowledge financial support from the Wellcome Trust IEH 102431, EPSRC (EP/S022104/1, EP/S013687/1), EPSRC Centre for Medical Engineering [WT 203148/Z/16/Z], the National Institute for Health Research (NIHR) Biomedical Research Centre (BRC) based at Guy’s and St Thomas’ NHS Foundation Trust and King’s College London and supported by the NIHR Clinical Research Facility (CRF) at Guy’s and St Thomas’. This work was also supported by the UK Research and Innovation London Medical Imaging and Artificial Intelligence Centre. We gratefully acknowledge the financial support provided by the German Academic Exchange Service (DAAD) through the Research Stipend (57745342), 2024.